A car from rest starts and accelerates 6 π‘š/𝑠 2 for 5 𝑠, after which it travels with a constant velocity for 9 𝑠. The brakes are then applied so that it decelerates at 4 π‘š/𝑠 2 . Find the total distance traveled by car.


1
Expert's answer
2021-09-24T09:27:07-0400

Given:

"v_1=0\:{\rm m/s},\quad a_1=6\:{\rm m/s^2},\quad t_1=5\:\rm s;"

"a_2=0\:{\rm m/s^2},\quad t_2=9\:\rm s;"

"a_3=-4\:{\rm m/s^2},\quad v_4=0\:\rm m/s."


The total distance

"d=d_1+d_2+d_3"

"d=(a_1t_1^2/2)+(a_1t_1*t_2)+\left(\frac{v_4^2-(a_1t_1)^2}{2a_3}\right)"

"=(6*5^2/2)+(6*5*9)+\left(\frac{0^2-(6*5)^2}{2*(-4)}\right)=457.5\:\rm m"